OpenStudy (anonymous):

Use the Identity of Property of multiplication, 33xy/3x, help?

OpenStudy (anonymous):

\[\frac{ 33xy }{ 3x }\]

OpenStudy (skullpatrol):

$$\huge \frac{ac}{bd}=\frac{a}{b}*\frac{c}{d}$$

OpenStudy (anonymous):

Here is a guess. Identity property of multiplication is simply when you multiply any number (or fraction or whatever you have) by 1 you get itself. Here we have \[\frac{ 33xy }{ 3x }=\frac{ 33yx }{ 3x }=\frac{ 33y }{ 3 }*\frac{ x }{ x }=\frac{ 33y }{ 3 }*1=\frac{ 33y }{ 3 }\] you simplify :-)

OpenStudy (anonymous):

The first step is using the fact that xy=yx since multiplication of numbers (real numbers) is commutative..example:. 3*4=4*3 the second step is using the fact that you can break up fractions and the third is that x/x=1

OpenStudy (anonymous):

@mashe THANK YOU!! IM SORRY FOR THE LATE RESPONSE)(T^T), THANK YOU SOO MUCH FOR THE HELP, YOU ARE AMAZING!!!

OpenStudy (anonymous):

you are welcome no prob

OpenStudy (anonymous):

Wait, @mashe I totally get everything you said and how you explained it was genius especially for a guess, but is there still more simplifying to do and I? (Btw, thank you again, I was stuck on this question for some minutes and I didn't think anyone would help, and then came you!^^)

OpenStudy (anonymous):

Yes I stated earlier you needed to simplify... you have \[\frac{ 33y }{ 3 }=\frac{ 33 }{ 3 }*\frac{ y }{ 1 }=11y\]

OpenStudy (anonymous):

OMG, I THINK LOVE YOU!!! THANK YOU SOO MUCHH !!!! <3333333 @mashe